# Heavy-Tailed Linear Bandits: Huber Regression with One-Pass Update

### v1

@@ 3 One-Pass Heavy-tailed Linear Bandits
In this section, we present our one-pass method for handling linear bandits with heavy-tailed noise. Section 3.1 introduces the problem setup. Section 3.2 reviews the latest work and identifies the inefficiency of the previous method. Section 3.3 introduces the OMD-type estimator based on Huber loss, along with the UCB-based arm selection strategy and the regret guarantee.
@@ 3.1 Problem Setup
We investigate the heavy-tailed linear bandits. At round $t$, the learner chooses an arm $X_{t}$ from the feasible set $\mathcal{X}_{t}\subseteq\mathbb{R}^{d}$, and then the environment reveals a reward $r_{t}\in\mathbb{R}$ such that
@@
We work under the following assumption as prior works (Xue et al., 2023; Huang et al., 2024).
@@ 3.2 Reviewing Previous Efforts
Efficiency Concern.   When the robustification parameter $\tau_{s}=+\infty$, Huber regression (6) reduces to least square, allowing one-pass updates with a closed-form solution. However, for finite $\tau_{s}$, which is essential for handling heavy-tailed noise, (6) requires solving a strongly convex and smooth optimization problem, which needs to store all past data resulting in a storage complexity of $\mathcal{O}(t)$. Using Gradient Descent (GD) to achieve an $\epsilon$-accurate solution requires $\mathcal{O}(\log(1/\epsilon))$ iterations, with per-iteration computational cost of $\mathcal{O}(t)$ to compute the gradient over all past data. To ensure an optimal regret, $\epsilon$ is typically set to $1/\sqrt{T}$, leading to a total computational cost of $\mathcal{O}(t\log T)$ at round $t$. This would be infeasible for a large-scale time horizon of $T$.
@@ 3.3 Hvt-UCB: Huber loss-based One-Pass Algorithm
The adaptive Huber regression method proposed by Huang et al. (2024) incurs significant computational burden in each round and requires storing all past data, which is impractical in real-world applications. Thus, we aim to design a one-pass algorithm for HvtLB based on the Huber loss (5).
@@
where $\sigma_{t}$ is the normalization factor and $\tau_{t}$ is the robustification parameter, both for controlling the impact of heavy-tailed noise.
To enable a one-pass update, instead of solving Huber regression (6) using all historical data, we will adopt the Online Mirror Descent (OMD) framework (Orabona, 2019). As a general and versatile framework for online optimization, OMD has been widely applied to various challenging adversarial online learning problems, such as online games (Rakhlin and Sridharan, 2013; Syrgkanis et al., 2015), adversarial bandits (Abernethy et al., 2008; Wei and Luo, 2018), and dynamic regret minimization (Jacobsen and Cutkosky, 2022; Zhao et al., 2024), among others. In contrast, we leverage its potential to address stochastic bandit problems, drawing inspiration from recent advancements in the study of logistic bandits (Zhang and Sugiyama, 2023). Specifically, we propose the following OMD-based update to estimate the unknown parameter,
@@
where $\Theta$ is the feasible set defined in Assumption 1 and $\mathcal{D}_{\psi_{t}}(\cdot,\cdot)$ is the Bregman divergence associated to the regularizer $\psi_{t}:\mathbb{R}^{d}\mapsto\mathbb{R}$. The choice of the regularizer is crucial because, in stochastic bandits, parameter estimation is subsequently used for confidence set construction. Therefore, we define the regularizer using the local information of the current round $t$ as
$\psi_{t}(\theta)=\frac{1}{2}\left\|\theta\right\|_{V_{t}}^{2},\mbox{ with }V_{t}\triangleq\lambda I+\frac{1}{\alpha}\sum_{s=1}^{t}\frac{X_{s}X_{s}^{\top}}{\sigma_{s}^{2}},V_{0}\triangleq\lambda I_{d}.$ (9)
The choice of this local-norm regularizer is both practical and compatible with the self-normalized concentration used to construct the Upper Confidence Bound (UCB) of the estimation error. Here, the coefficient $\alpha$ in $V_{t}$ essentially represents the step size of OMD, which will be specified later.
@@
$\widehat{\theta}_{t+1}=\widehat{\theta}_{t}-V_{t}^{-1}\nabla\ell_{t}(\widehat{\theta}_{t}),\quad\widehat{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\big{\|}\theta-\widehat{\theta}_{t+1}\big{\|}_{V_{t}},$
@@
$\begin{split}\sigma_{t}=\max\Big{\{}\nu_{t},\sigma_{\min},\sqrt{\frac{2\beta_{t-1}}{\tau_{0}\sqrt{\alpha}t^{\frac{1-\varepsilon}{2(1+\varepsilon)}}}}\left\|X_{t}\right\|_{V_{t-1}^{-1}}\Big{\}},\quad\tau_{t}=\tau_{0}\frac{\sqrt{1+w_{t}^{2}}}{w_{t}}t^{\frac{1-\varepsilon}{2(1+\varepsilon)}},\quad\tau_{0}=\frac{\sqrt{2\kappa}(\log 3T)^{\frac{1-\varepsilon}{2(1+\varepsilon)}}}{\left(\log\frac{2T^{2}}{\delta}\right)^{\frac{1}{1+\varepsilon}}},\end{split}$
@@
$\begin{split}\beta_{t}&\triangleq 107\log\frac{2T^{2}}{\delta}\tau_{0}t^{\frac{1-\varepsilon}{2(1+\varepsilon)}}+\sqrt{\lambda\left(2+4S^{2}\right)},\quad\text{where}~{}\kappa\triangleq d\log\left(1+\frac{L^{2}T}{4\sigma_{\min}^{2}\lambda d}\right).\end{split}$ (10)
The proof of Lemma 1 is provided in Appendix B.5. Notably, both of our method and the HeavyOFUL algorithm of Huang et al. (2024) require the input of $\sigma_{t},\tau_{t}$. In fact, our parameter setting for $\sigma_{t}$ is simpler. HeavyOFUL requires maintaining two complex lower bounds for $\sigma_{t}$ to satisfy: $\forall\theta\in\Theta,\left|\left(X_{t}^{\top}\theta-X_{t}^{\top}\theta_{*}\right)/\sigma_{t}\right|\leq\tau_{t}$ ensuring that the loss remains quadratic for noiseless data. In contrast, our analysis focuses on a refined condition to guarantee the quadratic loss property involving only the current estimate $\widehat{\theta}_{t}$: $\big{|}(X_{t}^{\top}\widehat{\theta}_{t}-X_{t}^{\top}\theta_{*})/\sigma_{t}\big{|}\leq\tau_{t}$. This refinement enables us to design a recursive parameter setting that directly incorporates the upper confidence bound $\beta_{t-1}$ from the previous round into the configuration of $\sigma_{t}$ for the current round. Furthermore, this parameter setting can be directly applied to the adaptive Huber regression. Similar to Huang et al. (2024), our method requires the knowledge of the moment $\nu_{t}$ of each round $t$. If $\nu_{t}$ is not available but a general bound $\nu$ is known, such that $\mathbb{E}[|\eta_{t}|^{1+\varepsilon}\mid\mathcal{F}_{t-1}]\leq\nu^{1+\varepsilon}$, we can still trivially achieve the same estimation error bound by replacing $\nu_{t}$ with $\nu$ in the setting of $\sigma_{t}$.
@@
Algorithm 1 Hvt-UCB
0:  time horizon $T$, confidence $\delta$, regularizer $\lambda$, $\sigma_{\min}$, bounded parameters $S$, $L$, parameter for estimation $c_{0}$, $c_{1}$, $\tau_{0}$, $\alpha$, $\varepsilon$.
2:  for $t=1,2,...,T$ do
3:     Select $X_{t}=\operatorname*{arg\,max}_{\mathbf{x}\in\mathcal{X}_{t}}\left\{\big{\langle}\mathbf{x},\widehat{\theta}_{t}\big{\rangle}+\beta_{t-1}\left\|\mathbf{x}\right\|_{V_{t-1}^{-1}}\right\}$
4:     Receive the reward $r_{t}$, and $\nu_{t}$
7:     Update $V_{t}=V_{t-1}+\alpha^{-1}\sigma_{t}^{-2}X_{t}X_{t}^{\top}$
8:     Compute $\widehat{\theta}_{t+1}$ by (8) and $\beta_{t}$ by (10)
9:  end for
@@ 4 Analysis Sketch
In this section, we provide a proof sketch for Lemma 1, where we adapt the analysis framework of Online Mirror Descent (OMD) to the SLB setting. To simplify the presentation, we first illustrate this framework using a simpler case where the noise follows a sub-Gaussian distribution in Section 4.1. We then discuss the challenges of extending this framework to heavy-tailed case and present our refined analysis in Section 4.2.
@@ 4.1 Case study of sub-Gaussian noise
$\begin{split}{}&\left\|\widehat{\theta}_{t+1}-\theta_{*}\right\|_{V_{t}}^{2}\leq\underbrace{2\sum_{s=1}^{t}\left\langle\nabla\widetilde{\ell}_{s}(\widehat{\theta}_{s})-\nabla\ell_{s}(\widehat{\theta}_{s}),\widehat{\theta}_{s}-\theta_{*}\right\rangle}_{\text{generalization gap term}}+\underbrace{\sum_{s=1}^{t}\left\|\nabla\ell_{s}(\widehat{\theta}_{s})\right\|^{2}_{V_{s}^{-1}}}_{\text{stability term}}-\underbrace{\frac{1}{2}\sum_{s=1}^{t}\left\|\widehat{\theta}_{s}-\theta_{*}\right\|_{\frac{X_{s}X_{s}^{\top}}{\sigma_{s}^{2}}}^{2}}_{\text{negative term}}+4\lambda S^{2},\end{split}$
@@
First based on the standard analysis of OMD (Orabona, 2019, Lemma 6.16), we can have the following inequality,
$\begin{split}\left\langle\nabla\ell_{t}(\widehat{\theta}_{t}),\widehat{\theta}_{t}-\theta_{*}\right\rangle\leq\frac{1}{2}\left\|\widehat{\theta}_{t}-\theta_{*}\right\|_{V_{t}}^{2}-\frac{1}{2}\left\|\widehat{\theta}_{t+1}-\theta_{*}\right\|_{V_{t}}^{2}+\frac{1}{2}\left\|\nabla\ell_{t}(\widehat{\theta}_{t})\right\|^{2}_{V_{t}^{-1}}.\end{split}$
@@ 4.2 Analysis for heavy-tailed noise
$\begin{split}\sum_{s=1}^{t}\left\|\nabla\ell_{s}(\widehat{\theta}_{s})\right\|^{2}_{V_{s}^{-1}}\leq\underbrace{2\sum_{s=1}^{t}\left(\min\left\{\left|\frac{\eta_{s}}{\sigma_{s}}\right|,\tau_{s}\right\}\right)^{2}\left\|\frac{X_{s}}{\sigma_{s}}\right\|^{2}_{V_{s}^{-1}}}_{\text{stochastic term}}+\underbrace{2\sum_{s=1}^{t}\left(\frac{X_{s}^{\top}\theta_{*}-X_{s}^{\top}\widehat{\theta}_{s}}{\sigma_{s}}\right)^{2}\left\|\frac{X_{s}}{\sigma_{s}}\right\|^{2}_{V_{s}^{-1}}}_{\text{deterministic term}}.\end{split}$
@@
$\begin{split}\underbrace{2\sum_{s=1}^{t}\left\langle\nabla\widetilde{\ell}_{s}(\widehat{\theta}_{s})+\nabla\ell_{s}(\theta_{*})-\nabla\ell_{s}(\widehat{\theta}_{s}),\widehat{\theta}_{s}-\theta_{*}\right\rangle}_{\text{Huber-loss term}}+\underbrace{2\sum_{s=1}^{t}\left\langle-\nabla\ell_{s}(\theta_{*}),\widehat{\theta}_{s}-\theta_{*}\right\rangle}_{\text{self-normalized term}},\end{split}$
@@
Finally, solving for $\beta_{t}$, we get $\beta_{t}=\mathcal{O}\left(t^{\frac{1-\varepsilon}{2(1+\varepsilon)}}\right)$. Thus we finish the proof of Lemma 1.

### v2

@@ 3 One-Pass Heavy-tailed Linear Bandits
In this section, we present our one-pass method for handling linear bandits with heavy-tailed noise. Section 3.1 introduces the problem setup. Section 3.2 reviews the latest work of Huang et al. (2023) and identifies the inefficiency of the previous method. Section 3.3 introduces the OMD-type estimator based on Huber loss, along with the UCB-based arm selection strategy and the regret guarantee.
@@ 3.1 Problem Setup
We investigate heavy-tailed linear bandits. At round $t$, the learner chooses an arm $X_{t}$ from the feasible set $\mathcal{X}_{t}\subseteq\mathbb{R}^{d}$, and then the environment reveals a reward $r_{t}\in\mathbb{R}$ such that
@@
We work under the following assumption as prior works (Xue et al., 2023; Huang et al., 2023).
@@ 3.2 Reviewing Previous Efforts
Remark 1.
Note that (Kang & Kim, 2023) and (Kang et al., 2024) also use Huber regression as the estimator in different bandit problems. Adaptive Huber regression differs from their methods in two key aspects: first, it uses a time-varying robustification parameter $\tau_{s}$; second, it introduces a normalization factor $\sigma_{s}$ to further control the loss. This normalization factor not only enables instance-dependent regret bounds but also ensures that the denoised data remains within the quadratic region of the Huber loss, such that the Hessian of the loss function is guaranteed to have a lower bound, which is essential for theoretical analysis. The two prior works based on Huber regression avoid such normalization but introduce other algorithmic components and assumptions to achieve a similar effect: (i) Kang & Kim (2023) incorporates a forced exploration mechanism along with a lower-bounded Hessian assumption; (ii) Kang et al. (2024) employs a local adaptive majorize-minimization technique together with a lower bounded covariance matrix assumption to guarantee a lower bounded Hessian.
Efficiency Concern.   When the robustification parameter $\tau_{s}=+\infty$, Huber regression (6) reduces to least square, allowing one-pass updates with a closed-form solution. However, for finite $\tau_{s}$, which is essential for handling heavy-tailed noise, (6) requires solving a strongly convex and smooth optimization problem, which needs to store all past data resulting in a storage complexity of $\mathcal{O}(t)$. Using Gradient Descent (GD) to achieve an $\epsilon$-accurate solution requires $\mathcal{O}(\log(1/\epsilon))$ iterations, with per-iteration computational cost of $\mathcal{O}(t)$ to compute the gradient over all past data. To ensure an optimal regret, $\epsilon$ is typically set to $1/\sqrt{T}$, leading to a total computational cost of $\mathcal{O}(t\log T)$ at round $t$. This would be infeasible for large-scale time horizon $T$.
@@ 3.3 Hvt-UCB: Huber loss-based One-Pass Algorithm
The adaptive Huber regression method proposed by Huang et al. (2023) incurs significant computational burden in each round and requires storing all past data, which is less favorable in practice. Thus, we aim to design a one-pass algorithm for HvtLB based on the Huber loss (5).
@@
where $\sigma_{t}$ is the normalization factor, and $\tau_{t}$ is the robustification parameter, both for controlling heavy-tailed noise.
To enable a one-pass update, instead of solving Huber regression (6) using all historical data, we adopt the Online Mirror Descent (OMD) framework (Orabona, 2019). As a general and versatile framework for online optimization, OMD has been widely applied to various challenging adversarial online learning problems, such as online games (Rakhlin & Sridharan, 2013; Syrgkanis et al., 2015), adversarial bandits (Abernethy et al., 2008; Wei & Luo, 2018), and dynamic regret minimization (Zhao et al., 2024; Jacobsen & Cutkosky, 2022), among others. In contrast, we leverage its potential to address stochastic bandit problems, drawing inspiration from recent advancements in the study of logistic bandits (Zhang & Sugiyama, 2023). Specifically, we propose the following OMD-based update for estimation,
@@
where $\Theta$ is the feasible set defined in Assumption 1 and $\mathcal{D}_{\psi_{t}}(\cdot,\cdot)$ is the Bregman divergence associated to the regularizer $\psi_{t}:\mathbb{R}^{d}\mapsto\mathbb{R}$. The choice of the regularizer is crucial because, in stochastic bandits, parameter estimation is subsequently used for confidence set construction. Therefore, we define the regularizer using the local information as
$\psi_{t}(\theta)=\frac{1}{2}\left\|\theta\right\|_{V_{t}}^{2},\mbox{ with }V_{t}\triangleq\lambda I+\frac{1}{\alpha}\sum_{s=1}^{t}\frac{X_{s}X_{s}^{\top}}{\sigma_{s}^{2}},$ (9)
where $V_{0}\triangleq\lambda I_{d}$. The choice of this local-norm regularizer is compatible with the self-normalized concentration used to construct the Upper Confidence Bound (UCB) of estimation error. Here, the coefficient $\alpha$ in $V_{t}$ essentially represents the step size of OMD, which will be specified later.
@@
$\widetilde{\theta}_{t+1}=\widehat{\theta}_{t}-V_{t}^{-1}\nabla\ell_{t}(\widehat{\theta}_{t}),\quad\widehat{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\big{\|}\theta-\widetilde{\theta}_{t+1}\big{\|}_{V_{t}},$
@@
$\begin{split}\sigma_{t}&=\max\Big{\{}\nu_{t},\sigma_{\min},\sqrt{\frac{2\beta_{t-1}}{\tau_{0}\sqrt{\alpha}t^{\frac{1-\varepsilon}{2(1+\varepsilon)}}}}\left\|X_{t}\right\|_{V_{t-1}^{-1}}\Big{\}},\\ \tau_{t}&=\tau_{0}\frac{\sqrt{1+w_{t}^{2}}}{w_{t}}t^{\frac{1-\varepsilon}{2(1+\varepsilon)}},\quad\tau_{0}=\frac{\sqrt{2\kappa}(\log 3T)^{\frac{1-\varepsilon}{2(1+\varepsilon)}}}{\left(\log\frac{2T^{2}}{\delta}\right)^{\frac{1}{1+\varepsilon}}},\end{split}$
@@
$\begin{split}\beta_{t}&\triangleq 107\log\frac{2T^{2}}{\delta}\tau_{0}t^{\frac{1-\varepsilon}{2(1+\varepsilon)}}+\sqrt{\lambda\left(2+4S^{2}\right)},\end{split}$ (10)
where $\kappa\triangleq d\log\left(1+\frac{L^{2}T}{4\sigma_{\min}^{2}\lambda d}\right)$.
We provide an analysis sketch in Section 4, where we explain how to analyze the estimation error using the OMD framework, and decompose the effects of heavy-tailed noise and one-pass updates. The full proof of Lemma 1 is provided in Appendix B.5.
Notably, both our method and the HeavyOFUL algorithm of Huang et al. (2023) require the input of $\sigma_{t},\tau_{t}$. In fact, our parameter setting for $\sigma_{t}$ is simpler. HeavyOFUL requires maintaining two complex lower bounds for $\sigma_{t}$ to satisfy: $\forall\theta\in\Theta,\left|\left(X_{t}^{\top}\theta-X_{t}^{\top}\theta_{*}\right)/\sigma_{t}\right|\leq\tau_{t}$ ensuring that the loss remains quadratic for noiseless data. In contrast, our analysis focuses on a refined condition to guarantee the quadratic loss property involving only the current estimate $\widehat{\theta}_{t}$: $\big{|}(X_{t}^{\top}\widehat{\theta}_{t}-X_{t}^{\top}\theta_{*})/\sigma_{t}\big{|}\leq\tau_{t}$. This refinement enables us to design a recursive parameter setting that directly incorporates the upper confidence bound $\beta_{t-1}$ from the previous round into the configuration of $\sigma_{t}$ for the current round. Furthermore, this parameter setting can be directly applied to the adaptive Huber regression. Similar to Huang et al. (2023), our method requires the knowledge of the moment $\nu_{t}$ of each round $t$. If $\nu_{t}$ is not available but a general bound $\nu$ is known, such that $\mathbb{E}[|\eta_{t}|^{1+\varepsilon}\mid\mathcal{F}_{t-1}]\leq\nu^{1+\varepsilon}$, we can still trivially achieve the same estimation error bound by replacing $\nu_{t}$ with $\nu$ in the setting of $\sigma_{t}$.
10:  time horizon $T$, confidence $\delta$, regularizer $\lambda$, $\sigma_{\min}$, bounded parameters $S$, $L$, parameter for estimation $c_{0}$, $c_{1}$, $\tau_{0}$, $\alpha$, $\varepsilon$.
2:  for $t=1,2,...,T$ do
3:     Select $X_{t}=\operatorname*{arg\,max}_{\mathbf{x}\in\mathcal{X}_{t}}\langle\mathbf{x},\widehat{\theta}_{t}\rangle+\beta_{t-1}\|X\|_{V_{t-1}^{-1}}$
4:     Receive the reward $r_{t}$, and $\nu_{t}$
7:     Update $V_{t}=V_{t-1}+\alpha^{-1}\sigma_{t}^{-2}X_{t}X_{t}^{\top}$
8:     Compute $\widehat{\theta}_{t+1}$ by (8) and $\beta_{t}$ by (10)
9:  end for
Algorithm 1 Hvt-UCB
@@ 4 Analysis Sketch
In this section, we provide a proof sketch for Lemma 1, where we adapt the analysis framework of Online Mirror Descent (OMD) to the SLB setting. To simplify the presentation, we first illustrate this framework using a simpler case where the noise follows a sub-Gaussian distribution in Section 4.1. We then discuss the challenges of extending this framework to handle heavy-tailed noise and present our refined analysis in Section 4.2.
@@ 4.1 Case study of sub-Gaussian noise
$\begin{split}{}&\left\|\widehat{\theta}_{t+1}-\theta_{*}\right\|_{V_{t}}^{2}\leq\underbrace{2\sum_{s=1}^{t}\left\langle\nabla\widetilde{\ell}_{s}(\widehat{\theta}_{s})-\nabla\ell_{s}(\widehat{\theta}_{s}),\widehat{\theta}_{s}-\theta_{*}\right\rangle}_{\text{generalization gap term}}\\ &+\underbrace{\sum_{s=1}^{t}\left\|\nabla\ell_{s}(\widehat{\theta}_{s})\right\|^{2}_{V_{s}^{-1}}}_{\text{stability term}}-\underbrace{\frac{1}{2}\sum_{s=1}^{t}\left\|\widehat{\theta}_{s}-\theta_{*}\right\|_{\frac{X_{s}X_{s}^{\top}}{\sigma_{s}^{2}}}^{2}}_{\text{negative term}}+4\lambda S^{2},\end{split}$
@@
First based on the standard analysis of OMD (Orabona, 2019, Lemma 6.16), we have
$\begin{split}{}&\left\langle\nabla\ell_{t}(\widehat{\theta}_{t}),\widehat{\theta}_{t}-\theta_{*}\right\rangle\\ \leq{}&\frac{1}{2}\left\|\widehat{\theta}_{t}-\theta_{*}\right\|_{V_{t}}^{2}-\frac{1}{2}\left\|\widehat{\theta}_{t+1}-\theta_{*}\right\|_{V_{t}}^{2}+\frac{1}{2}\left\|\nabla\ell_{t}(\widehat{\theta}_{t})\right\|^{2}_{V_{t}^{-1}}.\end{split}$
@@ 4.2 Analysis for heavy-tailed noise
$\begin{split}\sum_{s=1}^{t}\left\|\nabla\ell_{s}(\widehat{\theta}_{s})\right\|^{2}_{V_{s}^{-1}}&\leq\underbrace{2\sum_{s=1}^{t}\left(\min\left\{\left|\frac{\eta_{s}}{\sigma_{s}}\right|,\tau_{s}\right\}\right)^{2}\left\|\frac{X_{s}}{\sigma_{s}}\right\|^{2}_{V_{s}^{-1}}}_{\text{stochastic term}}\\ &+\underbrace{2\sum_{s=1}^{t}\left(\frac{X_{s}^{\top}\theta_{*}-X_{s}^{\top}\widehat{\theta}_{s}}{\sigma_{s}}\right)^{2}\left\|\frac{X_{s}}{\sigma_{s}}\right\|^{2}_{V_{s}^{-1}}}_{\text{deterministic term}}.\end{split}$
@@
$\begin{split}&\underbrace{2\sum_{s=1}^{t}\left\langle\nabla\widetilde{\ell}_{s}(\widehat{\theta}_{s})+\nabla\ell_{s}(\theta_{*})-\nabla\ell_{s}(\widehat{\theta}_{s}),\widehat{\theta}_{s}-\theta_{*}\right\rangle}_{\text{Huber-loss term}}\\ &\qquad\qquad+\underbrace{2\sum_{s=1}^{t}\left\langle-\nabla\ell_{s}(\theta_{*}),\widehat{\theta}_{s}-\theta_{*}\right\rangle}_{\text{self-normalized term}},\end{split}$
@@
Finally, solving for $\beta_{t}$, we get $\beta_{t}=\mathcal{O}(t^{\frac{1-\varepsilon}{2(1+\varepsilon)}})$. Thus we finish the proof of Lemma 1.
5 Experiments
In this section, we evaluate the empirical performance and time efficiency of our proposed Hvt-UCB algorithm. We present two experiments under heavy-tailed noise (Student’s $t$) and light-tailed noise (Gaussian). Furthermore, we provide additional experiments in Appendix E, including various heavy-tailed noises (Pareto, Lomax, Fisk), time-varying $\nu_{t}$, and changing arm sets $\mathcal{X}_{t}$, to further validate the effectiveness and robustness of our algorithm.
Settings.   We consider the linear model $r_{t}=X_{t}^{\top}\theta_{*}+\eta_{t}$, where the dimension $d=2$, the number of rounds $T=18000$, and the number of arms $n=50$. Each dimension of the feature vectors for the arms is uniformly sampled from $[-1,1]$ and subsequently rescaled to satisfy $L=1$. Similarly, $\theta_{*}$ is sampled in the same way and rescaled to satisfy $S=1$. We conduct two synthetic experiments with different distributions for noise $\eta_{t}$: (a) Student’s $t$-distribution with degree of freedom $df=2.1$ to represent heavy-tailed noise; and (b) Gaussian noise sampled from $\mathcal{N}(0,1)$ to represent light-tailed noise. For the heavy-tailed experiment, we set $\varepsilon=0.99$ and $\nu=1.31$, while for the light-tailed experiment, we set $\varepsilon=1$. We compare the performance of our proposed Hvt-UCB algorithm with the following baselines: (a) the OFUL algorithm (Abbasi-Yadkori et al., 2011); (b) the one-pass truncation-based algorithm CRTM (Xue et al., 2023); (c) the one-pass MOM-based algorithm CRMM (Xue et al., 2023); and (d) the Huber regression-based algorithm HEAVY-OFUL (Huang et al., 2023).
Figure 1: Student $t$: regret and time cost.
Figure 2: Gaussian: regret and time cost.
Results.   We conducted 10 independent trials and averaged the results. Figure 1 shows the cumulative regret and computational time under Student’s $t$ noise, while Figure 2 presents the same metrics under Gaussian noise, with shaded regions indicating the variance across trials. Under both noise settings, our algorithm demonstrates comparable regret performance to Heavy-OFUL while achieving remarkable computational efficiency, with a speedup exceeding 800×800$\times$800 ×. This efficiency advantage becomes more important as the number of rounds $T$ increases. OFUL, CTRM, and CRMM are faster than our Hvt-UCB algorithm because they use closed-form least squares solutions, while Hvt-UCB requires a projection step in each round. However, Hvt-UCB’s runtime remains competitive and is on the same order of magnitude as these least squares-based algorithms. Notably, among the one-pass algorithms, CRMM exhibits the fastest runtime because it updates its estimates periodically (every $\mathcal{O}(\log T)$ rounds) rather than in every round. In the heavy-tailed noise scenario (Figure 1), the performance of OFUL drops significantly, demonstrating its sensitivity to heavy-tailed distributions. In the Gaussian noise setting (Figure 2), our algorithm maintains competitive performance alongside OFUL. However, the truncation-based CRTM algorithm shows suboptimal performance due to excessive reward truncation in light-tailed noise scenarios, which leads to the loss of valuable data. These experimental results validate that our Hvt-UCB algorithm achieves robust performance across heavy-tailed and light-tailed noise environments while maintaining low computational cost.